OpenStudy (anonymous):

What is the expression in factored form? x^2 + 13x + 42

OpenStudy (anonymous):

(x+6)(x+7)

OpenStudy (anonymous):

How did you get that, @henryrodriguez713 ? ( http://openstudy.com/code-of-conduct )

OpenStudy (appleduardo):

6+7 = 13, 6*7 = 42 :P so henry is right, thats the factored form

OpenStudy (anonymous):

Sorry! Went to shower. Umm... This should help you with remembering the process of factoring quadratics. AND just know that if there is no "a" it's understood to represent 1 and if it's 1 you don't really have to divide, since 6/1 is 6. \[f(x)=ax^2 + bx +c\] MFDARM: My M:ultiply by a*c Father F:ind the factors of ac that add to b Drives D:ivide by a A This a <--- Red R:educe fractions Mustang M:ake into factors: put x on bottom and read up. \[f(x)=ax^2 + bx +c\] \[f(x)=x^2 + 13x + 42\] a=1 b=13 c=42 (Note: there is always an invisible 1 at "a") M: 1*42=42 F: 42 ** (NOTE: While getting the factors, add them and see if they add to 13)** 1 * 42 (1+42=43) NOT A FACTOR; doesn't add to 13 2 * 21 (2+42= 44) NOT A FACTOR; doesn't add to 13 3 * 14 (3+14=17) NOT A FACTOR; doesn't add to 13.( @thispointyouknowyou're close 6 * 7 (6+7=14) DING! DING! DING! Now divide by "a" \[\frac{ 6}{ 1x} \frac{ 7 }{ 1x }\] = (x+6)(x+7) So, (x+6)(x+7) and if you work it out you see it adds back to the original equation. NOTE: I know it doesn't make sense to divide by a, but it will help when a is 5 or 6 or 9 and so on.